martai:

Simplify the following expression. (7.3x+8.3)+(-4.14x+3.41

NoodlesAndRiceYT:

I see i see! Now to solve this, you need to use PEMDAS!

NoodlesAndRiceYT:

Now acorrding to PEMDAS, what do you do first?

martai:

idk

NoodlesAndRiceYT:

Well, do you know what PEMDAS is?

martai:

yes

martai:

Parentheses, Exponents, Multiplication and Division (from left to right), Addition and Subtraction (from left to right).

NoodlesAndRiceYT:

Yes! now according to all of that, what would you do first?

martai:

(7.3x+8.3)??

NoodlesAndRiceYT:

Ohh wait wait! So so sorry! i didnt see the x in the question! ahh, let me think.

martai:

oh nvm I got the answer

NoodlesAndRiceYT:

Oh okok, well sorry i got alittle confused and saw it wrong.

NoodlesAndRiceYT:

Glad you got the answer tho!

martai:

the answer is 11.44x+11.77

martai:

71

NoodlesAndRiceYT:

Since you got the answer, you can now close this question.

NoodlesAndRiceYT:

wait wait wait

martai:

?

NoodlesAndRiceYT:

How how did you get 71?

martai:

i mean 11.71

martai:

not 11.77

martai:

is 11.44x+11.71

NoodlesAndRiceYT:

OHH! okok sorry, i thought you meant 71 as in you multiplied, i was about to explain to you how that x isnt multiplication

NoodlesAndRiceYT:

Let me see if you are correct, i didnt check myself.

NoodlesAndRiceYT:

You are wrong!

martai:

how?

NoodlesAndRiceYT:

you got the 11.71 right, but not the other one.

NoodlesAndRiceYT:

look closely...

umm:

Given: (7.3x+8.3)+(-4.14x+3.41) =(7.3x+−4.14x)+(8.3+3.41) Step by step: Combine like terms =7.3x+8.3+−4.14x+3.41 =(7.3x+−4.14x)+(8.3+3.41) =3.16x+11.71 Therefore, your answer is 3.16x+11.71

NoodlesAndRiceYT:

7.3x and -4.14x

martai:

oh

NoodlesAndRiceYT:

put that together and you get...

NoodlesAndRiceYT:

oh ig someone else answered your question... ;-;

umm:

You can still answer it. Didn't mean to get in your way.

NoodlesAndRiceYT:

oh no its fine, i was just about done answering his question until you answered it, you did answer it good though, better then i did.

NoodlesAndRiceYT:

Since i got all confused in my way of answering the question but you didnt get confused and went right straight to it.

martai:

but thx for the help

umm:

No problem (: